OpenStudy (anonymous):

is anyone good at solving the system of linear equations by substitution help"??

OpenStudy (anonymous):

I can try what is the equation?

OpenStudy (anonymous):

I used to be good at this...

OpenStudy (anonymous):

it is 2x+y=5 y=x-4

OpenStudy (anonymous):

So it has already solved y for you, just plug in y in 2x+y=5 and solve for x then pick any of the following equations and plug x in to solve for y.

OpenStudy (anonymous):

2x+(x-4)=5

OpenStudy (anonymous):

okay so your equation is 2x +y = 5 ; y= x-4 my idea of this would to be plug i your information take the second part and plug it into the y of the first equation. 2x + (x-4) = 5 2x+x - 4 = 5 add like terms 2x and x then subtract the other terms tha tgo together -4 and 5 take the left over terms with variables and divide it by the terms that do not have variabls. if this is not clear please let me know.

OpenStudy (anonymous):

thanks to all of u :D